A 10 KG BLOCK LIES ON A FRICTIONLESS HORIZONTAL SURFACE.

A. WHAT FORCE WILL GIVE IT AN ACCELERATION OF 500 CM/S2 ?

B. WHAT CONSTANT FORCE WILL GIVE THE SAME BODY A VELOCITY OF 2 M/S AFTER 1 S ?


1
Expert's answer
2022-02-16T13:17:10-0500

Given:

"m=10\:\rm kg"

"a=500\:\rm cm/s^2=5\: m/s^2"

"\Delta v=2\:\rm m/s"

"\Delta t=1\: \rm s"


(a) the Newton's second law says

"F=ma=10*5=50\:\rm N"

(b)

"F=ma=m\frac{\Delta v}{\Delta t}=10*\frac{2}{1}=20\:\rm N"